OpenStudy (anonymous):

Suppose that the graph of f contains the point (3,6). Find a point that must be on the graph of g(x)=5(f(x)+2) This one's a little different.

hero (hero):

Not really

hero (hero):

g(3) = 5(f(3)+2)

hero (hero):

g(3) = 5(6+2) = 5(8) = 40

hero (hero):

3,40 is the point

hero (hero):

Check it

hero (hero):

Get it?

OpenStudy (anonymous):

ok im back. sorry.

OpenStudy (anonymous):

yeah, i think i get it. thanks.